The population of a region grows exponentially acoording to the function f(t)=6200(1.14)^t for t is less than or equal to 5 and greater than or equal to zero. Find the population of the region when t is 3 yeears.
a-10472
b-9186
c-9077
d-28272

Answer by stanbon(75887)

The population of a region grows exponentially acoording to the function f(t)=6200(1.14)^t for t is less than or equal to 5 and greater than or equal to zero. Find the population of the region when t is 3 yeears.
----------------
f(t)=6200(1.14)^t
f(3) = 6200(1.14)^4 = $9185.57
---------------------------------------
Cheers,
Stan H.
-----------------------
a-10472
b-9186
c-9077
d-28272